### v1

# Rethinking Visual Dependency in Long-Context Reasoning for Large Vision-Language Models

Abstract
Large Vision-Language Models (LVLMs) excel in cross-model tasks but experience performance declines in long-context reasoning due to overreliance on textual information and reduced visual dependency. In this study, we empirically analyze LVLMs in long-context reasoning, revealing that increased context length leads to a higher dependence on language at the expense of visual dependency. To address this issue, we propose a novel training-free context pruning method that selectively removes less critical textual information. Our approach enhances visual dependency and reduces textual noise, thereby improving LVLM performance in long-context reasoning. We validate our method by constructing a long-context dataset, demonstrating its effectiveness across various LVLMs. Moreover, further analysis confirms the robustness of different token pruning strategies and preliminary explores scaling laws between pruning rates and context length.
Rethinking Visual Dependency in Long-Context Reasoning
for Large Vision-Language Models
Yucheng Zhou1, Zhi Rao2, Jun Wan2, Jianbing Shen1thanks:  Corresponding author. 1 SKL-IOTSC, CIS, University of Macau 2 Macau University of Science and Technology yucheng.zhou@connect.um.edu.mo, jianbingshen@um.edu.mo
1 Introduction
Large Vision-Language Models (LVLMs) have achieved remarkable success across a range of cross-modal understanding and reasoning (Dai et al., 2023; Bai et al., 2023; Zhu et al., 2024). However, when LVLMs are applied to long dialogues or complex reasoning tasks, their performance tends to degrade. For instance, in our pilot experiments, LLaVA (Liu et al., 2023) exhibits up to a 17% decline in performance under long-context reasoning. This degradation shows that the model does not accurately understand the visual input, suggesting that LVLMs are overly dependent on textual priors, diminishing the visual dependency.
Recently, Large Language Models (LLMs) have significantly enhanced their performance in long-context reasoning through contextual position encoding and optimized attention strategies (Ding et al., 2024; Jin et al., 2024). These methods enable LLMs to retain and effectively utilize more information over long-context interactions, thereby maintaining contextual coherence and improving response accuracy. In contrast, LVLMs struggle to leverage visual information appropriately when processing long contexts. Consequently, it is essential for LVLMs to accurately integrate visual and textual information under long-context reasoning.
In this work, we conduct extensive analysis for LVLMs under long-context reasoning. Empirical results indicate a significant performance drop in LVLM as context length increases, while the opposite is true without visual information. It demonstrates that the model could rely more on the language priors with longer contexts, reducing its visual dependency. Moreover, our analysis further reveals that as the proportion of the target object within the image increases, the model’s visual dependency exhibits greater stability. By analyzing the attention weights within the model, we observe more cross-modal interaction in the shallow layers, followed by predominantly textual interaction in the deeper layers.
To improve the visual dependency in long-context reasoning, we propose a novel training-free context pruning method for LVLMs, which strategically reduces the interference of less critical textual information while preserving visual information. Our approach leverages the model’s internal attention mechanisms to identify and eliminate the trivial textual tokens based on their aggregated attention weights across multiple attention heads. By systematically pruning these less significant tokens, our method effectively diminishes textual noise and encourages more attention to visual information, enhancing visual dependency under long-context reasoning. In addition, we provide a theoretical foundation demonstrating that context pruning leads to more concentrated and stable attention distributions, which in turn fosters a stronger visual dependency on visual inputs.
To comprehensively evaluate the performance of LVLMs in long-context reasoning, we construct a long-context dataset based on the SVIT dataset (Zhao et al., 2023). Experimental results demonstrate the effectiveness of our approach across various LVLMs. Additionally, we further verify the robustness of our method by comparing different token pruning strategies, as well as various pruning layers and rates. The characteristics of our pruning method are further analyzed through part-of-speech observations of the pruned tokens. Lastly, we conduct a preliminary exploration of the scaling laws about the relationship between token pruning rates and context length.
In this study, our contributions are as below:
  • We conduct an empirical analysis of LVLMs in long-context reasoning, exploring the causes of performance decline and the factors influencing visual dependency.
  • We propose a novel train-free context pruning method that reduces less relevant textual information, improving visual dependency and performance in long-context reasoning.
  • We perform a theoretical analysis demonstrating that context pruning leads to more concentrated and stable attention distributions, which enhances visual dependency.
  • We construct a long-context dataset based on SVIT and validate our approach through comprehensive experiments, also analyzing pruning strategies and exploring scaling laws between pruning rates and context length.
2 Related Work
2.1 Large Vision-Language Models
Recent advancements in LVLMs have focused on enhancing cross-modal understanding by integrating visual and textual inputs for complex tasks (Wang et al., 2023; Chen et al., 2023a, b; Qu et al., 2024). For instance, LLaVA (Liu et al., 2023) leverages GPT-4 (OpenAI, 2023) to generate multimodal instruction-following data, improving performance in tasks like Science QA and multimodal reasoning. Expanding on this, models such as SVIT (Zhao et al., 2023) and LLaVA-NeXT (Liu et al., 2024a) utilize larger datasets and improve capabilities in OCR and abstract reasoning. Approaches like DeepSeek-VL (Lu et al., 2024) address real-world challenges by handling noisy and ambiguous inputs, while ViP-LLaVA (Cai et al., 2024) enhances robustness in processing diverse visual prompts. Additionally, some methods such as SURf’s self-refinement framework improve the accuracy of information retrieval (Sun et al., 2024), and VICL (Zhou et al., 2024) enables few-shot learning through inference on visual captioning. Zhang et al. (2024) extends to zero-shot video understanding using AnyRes, allowing models trained on images to generalize to videos, while Video-LLaMA2  (Cheng et al., 2024) enhances spatial-temporal modeling by integrating visual and audio inputs. These developments highlight LVLMs’ trends in scaling data, improving reasoning, and increasing robustness.
2.2 Long-Context Reasoning
Handling long contexts is a key challenge for LLM. Traditional position encoding techniques often struggle with long-range dependencies or varying-length sequences. To address this, methods like Contextual Position Encoding (CoPE) (Golovneva et al., 2024) dynamically adjust positional encodings based on context, improving performance on tasks requiring long-range dependencies. Training-free techniques, such as contextual sliding windows (An et al., 2024), extend the context length of pre-trained models without retraining, reducing computational overhead while maintaining sequence-processing capabilities. Parallel Context Windows (PCW) (Ratner et al., 2023) further enhances efficiency by splitting sequences into chunks for parallel processing. However, models tend to underutilize information from earlier inputs, favoring the middle or end (Liu et al., 2024b). Some methods like ThoT (Zhou et al., 2023) address this by improving reasoning across fragmented contexts, while LLMLingua (Jiang et al., 2023) compresses prompt representations, striking a balance between efficiency and context retention.
3 Background and Notation
LVLMs include a vision encoder and a LLM, and their input consists of images and text, which are processed through cross-modal interaction. The vision encoder maps the input image ${\bm{I}}$ to $n$ low-dimensional feature vectors ${\bm{V}}_{I}={\bm{v}}_{i}\in\mathbb{R}^{d},i\in\{1,2,\cdots,n\}$. While the text embedding layer transforms the input text sequence ${\bm{T}}=\{t_{1},t_{2},...,t_{m}\}$ into a sequence of embeddings ${\bm{H}}_{T}=\{{\bm{h}}_{1},{\bm{h}}_{2},...,{\bm{h}}_{m}\}$, where each ${\bm{h}}_{i}\in\mathbb{R}^{d}$ represents the embedding of the $i$-th token. The visual features ${\bm{V}}_{I}$ and the text embeddings ${\bm{H}}_{T}$ are fed into the LLM, predicting subsequent tokens in the text sequence by cross-modal interaction, i.e.,
$p(t_{m+1}|{\bm{T}},{\bm{I}})=\operatorname*{LLM}({\bm{H}}_{T},{\bm{V}}_{I}).$ (1)
The visual dependency ${\bm{D}}_{\text{vis}}(k)$ measures the extent to which the language model relies on visual features ${\bm{V}}_{I}$ when generating the $k$-th token. Specifically, the attention weights $\alpha_{k,i}$ at position $k$ are distributed among all available inputs, i.e., textual embeddings ${\bm{H}}_{T}$ and visual features ${\bm{V}}_{I}$. As the context length $m$ increases, the attention weights are normalized over a larger set of tokens, resulting in smaller individual attention weights for the visual features, i.e.,
$\mathcal{D}_{\text{vis}}(k)=\sum_{i=1}^{n}\alpha_{k,i},$ (2)
where $\alpha_{k,i}$ is the attention weight from position $k$ to visual feature vector ${\bm{v}}_{i}$. Since the attention mechanism normalizes the weights over all input tokens, increasing textual tokens $m$ can result in smaller attention weights $\alpha_{k,i}$ for the visual features. Consequently, the impact of ${\bm{V}}_{I}$ on the model’s predictions diminishes as the textual context grows, which presents a decrease in visual dependency.
4 Analysis on Visual Dependency of LVLMs in Long-Context Reasoning
4.1 Experimental Setting.
To analyze the visual dependency of LVLMs in long-context reasoning, we construct a long-context dataset based on the SVIT dataset (Zhao et al., 2023). SVIT dataset comprises images with multiple types of annotations, including conversation, complex reasoning, referring QAs, and detail description. These annotations can be concatenated to form long text annotations for images. We retain a total of 15,295 samples through random sampling and filtering. The samples are proportionally distributed into groups of 0.5k, 1k, 1.5k, 2k, and 2.5k based on text length. Each sample is processed in a question-answer format, with filtering to ensure that the objects in the answers do not appear in the context provided for the questions. For the analysis, we used the following LVLMs with 7B parameters: DeepSeek-VL (Lu et al., 2024), LLaVA (Liu et al., 2023), LLaVA-Next (Liu et al., 2024a), ViP-LLaVA (Cai et al., 2024), LLaVA-Next-Video (Zhang et al., 2024), and Video-LLaMA2 (Cheng et al., 2024).
4.2 Does Context Length Affect Visual Dependency?
Figure 1: Performance of various LVLMs on long-context reasoning.
To investigate whether context length affects the visual dependency of LVLMs, we conducted experiments on various LVLMs with progressively long contexts. As shown in Figure 1, we observe that the performance of various LVLMs declines as the context length increases. DeepSeek-VL experiences an approximate 28% drop in performance. In contrast, LVLMs trained on video data, i.e., LLaVA-NeXT-Video and Video-LLaMA2, exhibit greater robustness on long text, with relatively smaller reductions in performance. However, their performance also diminishes with increased context length. It indicates that as the text length increases, the dependency of LVLMs on visual input decreases, leading to a decline in performance.
Impact of Language Priors.
Figure 2: Performance of various LVLMs on text-only long contexts.
In addition to relying on visual input to infer the correct answers, LVLMs can also leverage language priors to make predictions. For instance, objects like “desk” and “chair” tend to co-occur frequently, allowing the model to use this co-occurrence probability as a cue. To investigate the impact of language priors, we replaced the images in our samples with either solid white or black images (to maintain identical token lengths) and fed these modified inputs into the LVLMs. As shown in Figure 2, most LVLMs perform better as the context length increases in the presence of strong language priors. This improvement occurs because longer contexts provide more details, making it easier for the model to infer the presence of related objects based on prior knowledge. However, LLaVA-Next maintains a relatively high performance overall. This can be attributed to the instruction tuning datasets of LLaVA-Next, which include same images from SVIT dataset. Consequently, LLaVA-Next has been trained to recognize and infer visual content based on contextual cues, even without images.
4.3 Consistency Between Vision and Language Priors
Figure 3: Consistency between vision and language priors.
To investigate the prediction consistency between vision and language priors, we analyze the disjoint portions of predictions with and without visual priors. Figure 3 presents the average performance of six models on different context lengths, comparing the accuracy of models provided visual priors (VL correct) and those only using language priors (Text correct). The proportion of samples that were correctly answered only when provided with an image generally decreases as the context length increases. It indicates that while visual dependency is more pronounced in shorter contexts, its impact diminishes as the model is exposed to more textual information. This suggests that the model increasingly relies on language reasoning under longer contexts. The proportion of samples only correctly answered without an image remains near zero across all context lengths. The results show that samples correctly answered using only language information remained to be correct even after adding visual information. It suggests that the models have effectively leveraged language priors to make accurate predictions, regardless of the availability of visual cues. As the context length increases, models rely more on the rich semantic information contained in the text, reducing visual dependency.
4.4 How Does Visual Content Affect Visual Dependency?
Figure 4: The relationship between the proportion of target visual content in the image and performance.
To analyze the impact of varying proportions of target visual content on model performance, we split samples based on the bounding box sizes of the target objects into three distinct groups: “Large”, “Medium” and “Small”. Figure 4 shows each category’s performance across different context lengths. From the figure, the model performance generally decreases as the context length increases. This trend is observed across all three groups. In particular, the group with “Large” bounding box values shows superior performance compared to the “Medium” and “Small” groups, which demonstrates that larger target objects (i.e., more significant visual content) allow the model to capture and utilize visual dependencies better.
4.5 Attention Weight Distribution Across Vision and Language Information
Figure 5: The distribution of attention weights between different layers of LVLMs.
To further investigate the model’s visual dependency, we analyze the distribution of attention weights in LVLMs from the perspective of cross-modal interactions. As shown in Figure 5, the attention distribution across different layers of the LLM backbone is visualized with 50 tokens per unit on the x-axis. The sequence is divided into two parts: the left side represents attention weights associated with visual information, and the right side corresponds to textual information. In the shallow layers (highlighted by the red box on the left), attention is heavily concentrated on the visual tokens, indicating that these layers are primarily responsible for processing and integrating visual inputs. Meanwhile, we observe a shift in attention towards the textual tokens (highlighted on the right), suggesting that deep layers focus more on language processing while still considering the visual context.
5 Context Pruning to Improve Visual Dependency
For long-context reasoning, the model’s attention weights become more distributed as the input length increases. It leads the model to rely more on textual priors, diminishing its dependence on visual content. As shown in Figure 6, we propose a training-free context pruning method designed to enhance the model’s visual dependency under long-context reasoning. By selectively pruning tokens in the textual input, we aim to reduce irrelevant information while retaining critical contextual cues. Moreover, we theoretically demonstrate the effectiveness and reliability of our method.
Figure 6: Overview of our method. Left: the token pruning process across multiple iterations; Right: pruning tokens in each LLM layer.
5.1 Context Pruning
Our method involves pruning a percentage $n\%$ of the least important tokens in the textual input based on their attention weights in the Transformer layers. Meanwhile, we ensure that visual tokens remain unpruned, allowing the model to retain complete access to visual information. The pruning is performed by first calculating the attention weights in each of the first $m$ layers of the Transformer and then ranking these weights to determine which tokens to prune. For tokens in multi-head attention, we perform max-pooling overheads to obtain a unified importance score for each token. We then prune $n\%$ tokens with the lowest attention scores.
Specifically, given a text $T=\{t_{1},t_{2},\cdots,t_{n}\}$ and an image $V=\{v_{1},v_{2},\cdots,v_{k}\}$, we compute the attention weights $A^{(l)}_{ij}$ between each pair of tokens $t_{i}$ and $t_{j}$ in $l$-th Transformer layer, i.e.,
$A^{(l)}_{ij}=\frac{\exp\left(\frac{Q^{(l)}_{i}\cdot K^{(l)}_{j}}{\sqrt{d_{k}}}\right)}{\sum_{k=1}^{n}\exp\left(\frac{Q^{(l)}_{i}\cdot K^{(l)}_{k}}{\sqrt{d_{k}}}\right)}$ (3)
where $Q^{(l)}_{i}$ and $K^{(l)}_{j}$ are the query and key representations of tokens $t_{i}$ and $t_{j}$, respectively, in layer $l$, and the attention weight $A^{(l)}_{ij}$ measures how much token $t_{i}$ attends to token $t_{j}$ in that layer. $\frac{1}{\sqrt{d_{k}}}$ is scaling factor, and $d_{k}$ the dimensionality of the key representations.
For multi-head attention, we aggregate the attention weights across all heads by applying max-pooling. The score for token $t_{i}$ is obtained by taking the maximum across all heads, i.e.,
$\hat{A}^{(l)}_{i}=\max_{h}\left(\sum_{j=1}^{n}A^{(l,h)}_{ij}\right)$ (4)
where $A^{(l,h)}_{ij}$ represent the attention weights for head $h$ in layer $l$. Once the scores $\hat{A}^{(l)}_{i}$ are calculated for all tokens, we sort these scores in ascending order and prune the bottom $n\%$ of tokens with the smallest scores. Only textual tokens $T$ are considered for pruning, while the visual tokens $V$ remain unpruned throughout the process. This ensures that visual information is fully preserved.
5.2 Theoretical Demonstration
Our method aims to improve LVLMs by selectively pruning text tokens, thereby reducing text noise and forcing the model to depend more on visual cues. We provide a theoretical demonstration to explain how pruning based on attention weights can lead to more stable and accurate attention distributions, and how this pruning enhances visual dependency.
Proposition 1:
After pruning low-importance tokens, the attention distribution becomes more concentrated on important tokens.
Proof:
Let $X=[x_{1},x_{2},\dots,x_{T}]$ be the input sequence of tokens at the $l$-th layer, where $T$ is the total number of tokens. The attention weight matrix $A^{(l)}$ contains the attention weights $A^{(l)}_{ij}$, where $A^{(l)}_{ij}$ denotes the attention weight of token $x_{i}$ over token $x_{j}$. The attention distribution for token $x_{i}$ is denoted as $A_{i}^{(l)}=[A^{(l)}_{i1},A^{(l)}_{i2},\cdots,A^{(l)}_{iT}]$, and can be seen as a probability distribution with $\sum_{j=1}^{T}A^{(l)}_{ij}=1$. The concentration of attention can be measured using entropy, i.e.,
$H(A_{i}^{(l)})=-\sum_{j=1}^{T}A^{(l)}_{ij}\log A^{(l)}_{ij}$ (5)
where smaller entropy indicates higher attention concentration, meaning the model focuses more on specific tokens. Consider pruning the lowest $n\%$ of tokens based on attention weights. Let the remaining tokens form the pruned sequence $\bar{X}=[\bar{x}_{1},\bar{x}_{2},\cdots,\bar{x}_{\bar{T}}]$, where $\bar{T}=(1-n\%)\times T$. After pruning, the attention distribution of token $x_{i}$ is updated to $\bar{A}^{(l)}_{i}$, and the entropy after pruning are defined as:
$H(\bar{A}^{(l)}_{i})=-\sum_{j=1}^{\bar{T}}\bar{A}^{(l)}_{ij}\log\bar{A}^{(l)}_{ij}$ (6)
Since the low-weight tokens have been pruned, the remaining attention weights are concentrated on more important tokens, leading to a reduction in entropy:
$H(\bar{A}^{(l)}_{i})\leq H(A^{(l)}_{i})$ (7)
Thus, pruning reduces the entropy of attention distributions, making the model’s focus more stable and accurate. Furthermore, to quantify the attention stability, we can calculate the variance of the attention weights before and after pruning for token $x_{i}$, i.e.,
$\sigma^{2}(A^{(l)}_{i})$ $=\frac{1}{T}\sum_{j=1}^{T}\left(A^{(l)}_{ij}-\mu(A^{(l)}_{i})\right)^{2}$
$\sigma^{2}(\bar{A}^{(l)}_{i})$ $=\frac{1}{\bar{T}}\sum_{j=1}^{\bar{T}}\left(\bar{A}^{(l)}_{i}-\mu(\bar{A}^{(l)}_{i})\right)^{2}$ (8)
where $\sigma^{2}(A^{(l)}_{i})$ and $\sigma^{2}(\bar{A}^{(l)}_{i})$ are the variance of the attention weights before and after pruning, respectively. Since pruning low-weight tokens increases the concentration of attention on the remaining tokens, the variance of the attention distribution increases:
$\sigma^{2}(\bar{A}^{(l)}_{i})\geq\sigma^{2}(A^{(l)}_{i})$ (9)
where the increase in variance indicates that the remaining attention is more focused on important tokens.
Proposition 2:
As textual tokens are pruned, the model becomes more dependent on visual information for reasoning.
Proof:
Consider a LVLM processes both visual input $V$ and textual input $T$ to produce an output $O$, such that $O=f(V,T)$. The model’s dependence on each modality can be quantified through its attention mechanism and the sensitivity of the output to each input. Let $A_{V}$ and $A_{T}$ represent the attention distributions over visual and textual inputs, respectively, satisfying the normalization condition:
$\sum_{i}A_{V,i}+\sum_{j}A_{T,j}=1,$ (10)
where $A_{V,i}$ and $A_{T,j}$ are the attention weights assigned to the $i$-th visual feature and the $j$-th textual token. Define $\alpha$ and $\beta$ as the total attention weights allocated to visual and textual inputs, respectively:
$\alpha=\sum_{i}A_{V,i},\quad\beta=\sum_{j}A_{T,j}.$ (11)
The model’s sensitivity to each modality can be linked to these attention weights. Specifically, the sensitivity of the output $O$ to the visual input $V$ is proportional to $\alpha$, and the sensitivity to the textual input $T$ is proportional to $\beta$:
$\left\|\frac{\partial O}{\partial V}\right\|\propto\alpha,\quad\left\|\frac{\partial O}{\partial T}\right\|\propto\beta.$ (12)
When textual tokens are pruned, the textual input reduces from $T$ to a pruned version $\bar{T}$, containing fewer tokens. Consequently, the attention weights over textual inputs decrease, leading to a reduced $\beta$:
$\beta^{\prime}=\sum_{j}\bar{A}_{T,j}<\beta,$ (13)
where $\bar{A}_{T,j}$ are the attention weights assigned to the remaining textual tokens after pruning. Since the total attention must still sum to 1, the attention allocated to visual inputs increases:
$\alpha^{\prime}=\sum_{i}\bar{A}_{V,i}=1-\beta^{\prime}>1-\beta=\alpha.$ (14)
where $\alpha^{\prime}>\alpha$ indicates increased dependence on visual inputs. This increase in $\alpha$ directly translates to greater sensitivity of the output $O$ to the visual input $V$:
$\left\|\frac{\partial O}{\partial V}\right\|^{\prime}\propto\alpha^{\prime}>\alpha\propto\left\|\frac{\partial O}{\partial V}\right\|.$ (15)
Similarly, the sensitivity to the textual input decreases:
$\left\|\frac{\partial O}{\partial\bar{T}}\right\|^{\prime}\propto\beta^{\prime}<\beta\propto\left\|\frac{\partial O}{\partial T}\right\|.$ (16)
Therefore, as textual tokens are pruned, the model reallocates its attention from textual to visual inputs, resulting in increased $\alpha^{\prime}$ and decreased $\beta^{\prime}$. This shift enhances the model’s dependency on visual information for reasoning, as evidenced by both the attention weights and the sensitivity measures.
6 Experiments
6.1 Experimental Setting
To validate the effectiveness of our proposed method, we conduct experiments using the long-context dataset collected as mentioned in Sec. 4.1, which is based on the SVIT dataset (Zhao et al., 2023). It comprises 15,295 samples, divided into five groups with 0.5k, 1k, 1.5k, 2k, and 2.5k samples, respectively. The LVLMs with 7B parameters used in our experiments include DeepSeek-VL, LLaVA, LLaVA-Next, ViP-LLaVA, LLaVA-Next-Video, and Video-LLaMA2. The token pruning is applied to the first 10 layers, with a pruning rate of 0.3. For evaluation, we utilize greedy decoding for LVLMs and conduct model inference on NVIDIA A100 GPUs.
6.2 Results
Figure 7: Top: Performance comparison of our method and various LVLMs. Bottom: Performance gain of our method compared with baselines.
In Figure 7, the performance of different LVLMs is compared across varying context lengths, with the top figure showing the performance of various LVLMs with our method and the bottom figure showing the performance gain of our method compared with baselines. Our method demonstrates strong and consistent performance, especially with Video-LaMA2, which maintains high performance across contexts of different lengths, indicating robustness to different context lengths. Although most models show a slight drop in performance as the context length increases, the performance gain analysis reveals that our method is especially effective in improving models that struggle with handling long contexts, with more significant gains observed for models with lower baseline performance. Even if models that already perform well at baseline show improvements, suggesting our method can provide refinement for LVLMs under long-context reasoning. In addition, due to token pruning, our method also offers a significant advantage in inference efficiency, reducing inference time to approximately half of the original.
6.3 Analysis on Token Pruning
Impact of Token Pruning Method.
Figure 8: Results with different token pruning methods.
Figure 8 shows the performance comparison of different token pruning methods: “Max Pooling”, “Mean Pooling”, and “Random”. Both “Max Pooling” and “Mean Pooling” maintain high accuracy as the sequence length decreases, supporting our claim in Proposition 1 that pruning low-importance tokens is crucial. These methods leverage multi-head attention to preserve key tokens with higher attention weights, ensuring minimal performance loss. In contrast, “Random” leads to a complete loss in accuracy, highlighting the importance of intelligently selecting tokens for pruning to maintain model performance.
Differences in Token Pruning Layers.
Figure 9: Results with different token pruning layers.
To investigate the impact of token pruning at different layers, we analyze the performance of our method when pruning tokens at the shallow, intermediate, and deep layers for varying input lengths. As shown in Figure 9, results demonstrate that token pruning in the shallow layers consistently achieves better performance across all tested input lengths, while pruning in later layers (i.e., intermediate and deep layers) results in a noticeable performance drop. The performance remains relatively unchanged compared to the baseline when pruning is conducted in the later layers. These findings confirm our observations from Figure 5, where we noted that attention weights between text and vision interactions are more significant in the shallow layers of LVLMs. This indicates that the shallow layers are more effective at identifying tokens that are less relevant to the visual information, making them ideal for filtering. By pruning tokens in these layers, we can preserve the most important information for cross-modal understanding.
Impact of Token Pruning layer Numbers.
Figure 10: Results with different token pruning layer numbers.
Figure 10 shows the impact of applying Token Pruning at different layer depths (7, 10, and 13 layers) across varying sequence lengths. The results indicate that deeper layers (10 and 13) generally achieve higher accuracy, particularly for longer sequences. While the accuracy difference is minimal for shorter sequences (0.5k and 1.0k tokens), it becomes more pronounced as the sequence length increases, with 10 layers outperforming other layers significantly at 2.5k tokens.
Impact of Token Pruning Rate.
Figure 11: Results with different token pruning rates.
To analyze the impact of token pruning rates on performance, we evaluate our method with different pruning rates (0.1, 0.2, and 0.3) across varying input lengths (0.5k to 2.5k tokens), as shown in Figure 11. The results indicate that for shorter sequences (0.5k and 1.0k), performance remains high across all pruning rates, with only a slight drop as the pruning rate increases. However, for longer sequences (1.5k to 2.5k), higher pruning rates lead to an improvement in performance, especially at the 0.3 pruning rate compared to 0.1 and 0.2. It demonstrates that a fixed pruning rate may not be suitable for all input lengths. Therefore, a dynamic pruning strategy that adjusts the rate based on the input length could better preserve accuracy across different sequence sizes.
Pruned Token Distribution.
Figure 12: The remaining tokens after pruning across different layers of the model.
Figure 12 shows the distribution of token types remaining after pruning at each layer of the model. It can be observed that although the proportion of remaining token types is relatively uniform across layers, different parts of speech exhibit varying degrees of retention as pruning progresses. For instance, nouns and verbs consistently make up a significant portion of the retained tokens, while prepositions and tokens in the “other” category show greater variability across layers. This result is in stark contrast to the performance degradation observed in Figure 8, where random token pruning leads to significantly worse performance. Although both approaches maintain a relatively uniform token pruning ratio across layers, the random pruning method fails to preserve critical information, as evidenced by its poor performance. However, results from Figure 12 demonstrate that our pruning method effectively identifies and removes less important tokens, retaining those that are more crucial for model performance. This highlights the strength of our approach in accurately selecting the tokens to prune, thus preserving model performance even under significant pruning.
6.4 Preliminary Exploration on Token Pruning Scaling Laws
Figure 13: Token Pruning Scaling Laws.
Figure 14: Performance of LVLMs with different token pruning rates and context lengths.
Figure 15: Case of token pruning using our method, with the red parts indicating the retained tokens.
Figure 13 shows the relationship between the pruning rate and performance in long-context reasoning. The “best pruning rate” represents averaging the top three adjacent performance points from testing various pruning rates. The performance for different pruning rates and lengths is shown in Figure 14. The plotted points correspond to the best pruning rates at different context lengths, while the red curve represents a fitted quadratic regression line, capturing the trend in the relationship between these variables. From the curve and data points, we observe a clear upward trend, indicating that as the token delete rate increases (i.e., as the context length extends), the best delete rate also rises. It demonstrates that for longer contexts, more aggressive pruning may be advantageous in preserving optimal performance.
6.5 Case Study
Our method effectively removes most irrelevant tokens, retaining only those related to the context. As shown in Fig. 15, the baseline method consistently fails in long-context reasoning tasks with contexts longer than 0.5k tokens, whereas our approach correctly answers across various context lengths.
7 Conclusion
In this study, we investigate the challenges faced by LVLMs in long-context reasoning. Our empirical analysis demonstrated that LVLMs experience a significant decline in visual dependency as the context length increases. We also identified that a higher proportion of target objects within images helps stabilize visual dependency and that cross-modal interactions are more prominent in the shallow layers of LVLMs, shifting towards predominantly textual interactions in deeper layers. To mitigate the performance degradation, we proposed a novel context pruning method that strategically removes less critical textual information. Our approach was validated using a newly constructed long-context dataset based on SVIT, and it proved effective across various LVLMs. Additional experiments confirmed the robustness of different token pruning strategies and elucidated the relationship between pruning rates and context length.
Limitations
While this study provides significant insights into the performance of LVLMs in long-context reasoning, there is still room for further research. The extensive computational resources required for our analysis limited the number of LVLMs we were able to evaluate. Additionally, our investigation into Token Pruning Scaling Laws was conducted as a preliminary exploration due to resource constraints. Although we identified a relationship between context length and Token Pruning rate, more comprehensive studies are needed to fully elucidate this connection and to establish robust scaling laws.